An open U tube of uniform cross section contains mercury .If 27.2 cm of water column is poured into one climb , how high does the mercury surface rise in other climb from its initial volume.?

Grade:12th pass

3 Answers

Abcd
6 years ago
P○ + density of water × height of water ×g= P○ + density of mercury ×height of mercury [ raised height] ×g
1×27.2=13.6 ×h
h= 2cm
The level pf mercury rise by 2 cm from its initial volume
Abcd
6 years ago
When the water is poured into one limb the mercury level in that limb decreases by h height and in creases in the other by h
Therefore pressure of decreased hieght h and the same level in other limb is sme 
And hence the height for the second limb should be taken as 2h
P○ + 27.2×1×g=13.6×2h×g
27.2=13.6×2h
2h=2h=1
 
ankit singh
3 years ago
When the water is poured into one limb the mercury level in that limb decreases by h height and in creases in the other by h
Therefore pressure of decreased hieght h and the same level in other limb is sme
And hence the height for the second limb should be taken as 2h
P○ + 27.2×1×g=13.6×2h×g
27.2=13.6×2h
2h=2h=1